OpenStudy (starryknight):

Find the sum of the first 33 terms of the following sequence: 8, 17, 26 . . . answers: 608 2376 5016 10032

hartnn (hartnn):

can you find the common difference between the terms ?

hartnn (hartnn):

or know what common difference is ?

OpenStudy (starryknight):

-9

hartnn (hartnn):

-9 or +9 ?

hartnn (hartnn):

d= 2nd term - 1st term = 3rd term - 2nd term = .... and so on

hartnn (hartnn):

d = 26- 17 = 17- 8 = ..?

OpenStudy (starryknight):

9

OpenStudy (starryknight):

17-8=9

hartnn (hartnn):

yes

hartnn (hartnn):

now apply the formula \(\Large S_n = (n/2) (2a+(n-1)d)\)

hartnn (hartnn):

where a =1st term = 8 n = number of terms = 33 d= 9

OpenStudy (starryknight):

s33=33/2(2*8(33-1)9) 33/2(2*8(32)9) 33/2(2*8(288) 33/2(16(288) 33/2(4608)

OpenStudy (starryknight):

something is wong i dont know what though

hartnn (hartnn):

there is a plus sign in between 2*8 and (33-1)9

hartnn (hartnn):

33/2(2*8+ (33-1)9)

OpenStudy (starryknight):

ooooh ok

OpenStudy (starryknight):

i got 5016

OpenStudy (starryknight):

thanx for sticking around to help me ^-^

hartnn (hartnn):

thats correct :) and welcome ^_^